A triangle has a perimeter of 57 inches. Side A is 3 inches shorter than twice side B, and side C is 4 inches longer than side B. How many inches long is side A?
Thanks for reading, I hope you answer soon.

Answer by stanbon(75887)

A triangle has a perimeter of 57 inches. Side A is 3 inches shorter than twice side B, and side C is 4 inches longer than side B. How many inches long is side A?
---------
Equations:
A + B + C = 57 inches
A = 2B-3
C = B+4
=========
Substitute for A and C and solve for "B":
2B-3 + B + B+4 = 57
------
4B + 1 = 57
4B = 56
B = 14 inches
----
A = 2B-3 = 2*14-3 = 25 inches
C = B+4 = 18 inches
-----------------------
Cheers,
Stan H.
-----------------------